Grade 11Trigonometry

tanx +tan4x +tan7x=tanxtan4xtan7x

pawan punasya
15 Years agoGrade 11

2 Answers

Akshith Yeso
14 Years ago

Given  tanx+tan4x+tan7x=tanxtan4tan7x

We know that tan(a+b+c)=tana+tanb+tanc-tanatanbtanc/1-(tanatanb+tanbtanc+tanctana)

           If a+b+c is n(pie),n belongs to N then tan(a+b+c)=0

                           => tana+tanb+tanc-tanatanbtanc=0

                           =>tana+tanb+tanc=tanatanbtanc

           it is same as the given question,

                        Hence , x+4x+7x=n(pie)

                                 =>12x=n(pie)

                                 => x=n(pie)/12   

ruchi yadav
12 Years ago

Given tanx+tan4x+tan7x=tanxtan4tan7x


We know that tan(a+b+c)=tana+tanb+tanc-tanatanbtanc/1-(tanatanb+tanbtanc+tanctana)


If a+b+c is n(pie),n belongs to N then tan(a+b+c)=0


=> tana+tanb+tanc-tanatanbtanc=0


=>tana+tanb+tanc=tanatanbtanc


it is same as the given question,


Hence , x+4x+7x=n(pie)


=>12x=n(pie)


=>x=n(pie)/12


Thank You
Ruchi
Askiitians Faculty